OpenStudy (haleyelizabeth2017):

Find the amplitude and period of f(x)=-5sin(5x-2).

Nnesha (nnesha):

\[\large\rm y=Asin (Bx-C )+D \] original equation where \[\left| A \right|\] is amplitude and \[ \frac{ 2\pi }{ \left| B \right| }\] period and D is vertical shift C is horizontal or phase shift

OpenStudy (haleyelizabeth2017):

Okay so 5 is the amp and 2pi/5 is period.

OpenStudy (haleyelizabeth2017):

How do I find this for a cosine function?

Nnesha (nnesha):

2pi is full cycle if given is degree then u should put 360 and yes that's correct

Nnesha (nnesha):

same formula for cosine function

OpenStudy (haleyelizabeth2017):

Okay, thank you very much :)